How much heat must be added to a system at 298 K for the number of accessible states to increase by a factor 


1
Expert's answer
2021-12-14T15:08:37-0500

How much heat must be added to a system at 298 K for the number of accessible states to increase by a factor of 10^6 ?

"Q = kT ln(W)"

where k is constant "= 1.38 \times 10^{-23 }"

Heat

"Q = (1.38 \times 10^{-23}) \times 298 \times ln (10^6) \\

Q = 5.68 \times 10^{-20} \; J \\

Q = 0.3544 \; eV"